OpenStudy (anonymous):

1 1/6 x2/5 someone solve .Remember to do in simplest form

OpenStudy (turingtest):

(7/6)(2/5)=14/30=7/15

OpenStudy (anonymous):

your really on it great excellent

OpenStudy (turingtest):

medals are better than thx, but you're welcome